A 340 kg marble boulder rolls off the top of a cliff and falls a vertical height of 160 m before striking the ground.

part A: Estimate the temperature rise of the rock if 42 % of the heat generated remains in the rock. Express your answer using two significant figures.
Answer #1

Given,

The mass of a marble, m = 340 Kg

The vertical height, y = 160 m

The Kinetic energy at the bottom is obtained from,

K = U = m g h

= m g y

= 340 * 9.8 * 160 = 5.33 * 105 J

42 % of K.E = 0.42 * 5.33 * 105

= 2.34 * 105 J

E = m c T

2.34 * 105 = 340 * 880 * T

T = 0.782 oC